OpenStudy (anonymous):

help please! solve for x:3(x+1)=-2(x-1)+6

OpenStudy (anonymous):

ok so youre going to want to begin by distributing the 3 into the (x+1) and the -2 into the (x-1)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so you should have 3x+3=-2x+1+6, right?

OpenStudy (anonymous):

yeah sorry my laptop was going nuts it wouldn't let me type!

OpenStudy (anonymous):

MY laptop isint working so if I don't reply just keep going I'm still here.

OpenStudy (anonymous):

haha all good all good! :) ok so next you want to combine all the x terms and all the terms without x: 3x+3=-2x+7, getting all the x's to one side and non x terms to the other yields: 3x+2x=7-3 which simplifies to 5x=4 or x=5/4

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yep!

OpenStudy (johnn.):

Let's solve your equation step-by-step. 3(x+1)=−2(x−1)+6 Step 1: Simplify both sides of the equation. 3(x+1)=−2(x−1)+6 Simplify: 3x+3=−2x+8 Step 2: Add 2x to both sides. 3x+3+2x=−2x+8+2x 5x+3=8 Step 3: Subtract 3 from both sides. 5x+3−3=8−3 5x=5 Step 4: Divide both sides by 5. 5x5=55 x=1 Answer: x=1

OpenStudy (anonymous):

^^hes correct, i must have made a goof somewhere! i apologize!